OpenStudy (anonymous):

Find the area of the triangle with vertices (1,1,5) (3,4,3) and (1,5,7)

OpenStudy (anonymous):

I got 8.307 am I right?

OpenStudy (zzr0ck3r):

what is the perimeter?

OpenStudy (ybarrap):

For (Xa,Xb,Xc)=(1,1,5),(Ya,Yb,Yc)= (3,4,3)and (Za,Zb,Zc)=(1,5,7): http://upload.wikimedia.org/math/1/7/c/17cb8db9ccd42b47a704ba683c7c5f5b.png Where | | is the determinant.

OpenStudy (ybarrap):

Correction: For (Xa,Ya,Za)=(1,1,5),(Xb,Yb,Zb)= (3,4,3)and (Xc,Yc,Zc)=(1,5,7): \(\frac{1}{2}\sqrt{(-4)^{2}+8^{2}+14^{2}}=\sqrt{69} \approx \) 8.307

OpenStudy (zale101):

Area of triangle = 1/2 * |axb| Let a be the vector from (1,1,5) to (3,4,3) Let b be the vector from (3,4,3) to (1,5,7) vector a = (1,1,5) - (3,4,3) = (-2,-3,2) vector b = (3,4,3) - (1,5,7)= (2,-1,-4) After that, find the cross product a x b | i j k | | -2i,-3j,2k| | 2i,-1j,-4k| =(14i,-4j,8k) since you’ve got the answer for the cross product now find the area of parallelogram of vector b and a. |a x b| = sqrt( (14)^2 + (-4)^2 + (8)^2)= sqrt 69 area of triangle = 1/2 * sqrt (69)=8.307! Thanks @ybarrap , i did something wrong with my calculations :)

OpenStudy (ybarrap):

Your solution is a lot more elegant. Amazing the number of ways to find areas!